OpenStudy (anonymous):

Please help! Im in such a hurry to get this done! medals rewarded. these are the statement, please give the reasons to these statments. the multiple choices are below. Here are the statments: 1) 5(x + 3) + 9 = 59 2) 5x + 15 + 9 = 59 3) 5x + 24 = 59 4) 5x + 0 = 35 5) 5x = 35 6) x = 7 now these are the Reason: for each statement, please give the reason, for the six of these statments; A. Reflexive Property B. Symmetric Property C. Transitive Property D. Addition Property of Equality E. Subtraction Property of Equality F. Multiplication Property of Equality G. Division Property of Equality H. Substitution Property I. Additive Identity J. Mult

OpenStudy (anonymous):

J. Multiplicative Identity K. Commutative Property L. Associative Property M. Distributive Property N. Given

hero (hero):

What is preventing you from doing this on your own?

OpenStudy (anonymous):

I have so many assignment to get caught up on right now. and I have no time at all. and it would be great if someone could help me. Plus im not really sure how to do this.

hero (hero):

Why did you wait until the last minute?

OpenStudy (anonymous):

i didnt. please just understand where I am comming from. I did not wait till the last minute. I take online school. and you can get behind very quick. I have to do 20 assignments a day. its a lot harder than people thingk. all im trying to do is get some help.

hero (hero):

20 assignments a day? You might be exaggerating a little.

OpenStudy (anonymous):

no I have to do 5 assignments for each class a day, and I have six classes. and have to spend one hour doing each class./

OpenStudy (anonymous):

1)5(x + 3) + 9 = 59 given prob? yes or no? give it a little thinking..lol :D 2) 5(x + 3)=5x+15 Multiplication ? 3)5x + (15 + 9) = 59 5x + 24 = 59 thats addition ? 4) 5x + 24 = 59 5x + 0 = 35 subtraction ? 5) 5x = 35 di you add 0? 6) x = 7 division? good luck now :D

OpenStudy (anonymous):

wouldnt the first one be a given

OpenStudy (anonymous):

and number 5 would be addition? @mark_o.

OpenStudy (anonymous):

i need the whole reason

OpenStudy (anonymous):

yes thats add 0